# morphisms_everywhere.bib

### v1

@@ 1. Introduction
Finding compatible operations in two domains is a highly developed essential tool in mathematics, but in fact, it is a universal mechanism for any understanding, reasoning and explanation, for human as well as for artificial intelligence.
These mechanisms are called morphisms (i.e. structure-preserving maps), and they could and should be used correctly and explicitly in all areas of knowledge to make philosophical problems more precise and reduce the chance of confusion arising from ambiguities.
@@
In Section 2 we go through some explicit and implicit uses of morphisms. To avoid further misunderstandings, we precisely define morphisms and fix a notation in Section 3. We discuss the canonical examples in Section 4. Section 5 elaborates on how the idea of morphisms underlie various aspects of fields related to intelligence, like explanation and understanding. We conclude in Section 6 with a summary and a research plan.
2. Background and Related Works
2.1. Category Theory
Informally, category theory can be described as an abstract theory of functions ([awodey2010category]). As a relatively concrete example, we mention algebraic objects of the same kind (e.g., groups, rings). The arrows between the objects represent homomorphisms. Visually, a category is a directed graph, where associative composition is defined. For a path, a composable sequence of arrows, there is an arrow for the corresponding composition. In other words, whenever there is a way (no matter how long) from one object to another, there has to be a direct link. It is rather surprising this structure can serve as the model of numerous mathematical models ([spivak2017]).
  1. (1)
    too specific to mathematics;
  2. (2)
    too general about morphisms.
There is no contradiction here. Learning category theory requires wide and mature mathematical knowledge, although recent books ([cheng2022joy, 2022_sheaf_Rosiak]) ease this situation. Also, the exact definition of a category is geared towards algebraic objects. The need for the identity morphism for each object is less relevant for a general philosophical theory of explanation and representation. On the other hand, a morphism can be anything in a category. For instance, a partially ordered set is a category, where the arrows, the morphisms are the $\leq$ relations. These arrows merely indicate the presence of the relation without being morphisms in the literal sense.
The most relevant feature of category theory here is that it provides tools from transforming one type of mathematics into another type. Moreover, category theory has a self-similar nature: a functor is a structure preserving map between categories, and natural transformation are morphisms between functors. Thus, we have a category of categories (assuming provisions against the paradoxes of naive set theory). This plurality of cross-level structure preserving maps is what we envision here for a wider context of the philosophical discourse about the interrelated notions of scientific explanation, understanding, reasoning, representation, and intelligence in general.
2.2. Algebraic Automata Theory
Algebraic automata theory provides a mathematical theory for (scientific) understanding. One of its fundamental result is the prime decomposition theorem of finite state automata ([primedecomp65]). It states that any automaton can be built from simpler components in a hierarchical way using series-parallel connections. Such a decomposition introduces coordinates into a discrete dynamical system ([wildbook]). These coordinates are not independent. Control information is unidirectional between the components, thus we have hierarchical coordinate systems. By omitting lower level coordinates we get a coarse-grained description of the system. For instance, conserved quantities in physical systems (e.g., energy) give the top level coordinates in a hierarchical decomposition.
Understanding in this framework is defined by giving a surjective morphism (emulation) from the explanandum into a nicer, easy to understand hierarchical structure. As we can understand something in several ways, the decompositions are not unique. For our discussion, the idea of different morphisms used for different purposes is the most relevant.
2.3. Cognitive Metaphors
Cognitive metaphors ([lakoff2008metaphors]) are thought processes where we “experience and understand one kind of things in terms of another”. Metaphor is a map between a source domain (what we know) and the target domain (what is new for us) with the purpose of conceptualizing the target. This is exactly the setup for a morphism, but the correspondence between the domains is not detailed. The authors indeed started with the idea of a mathematical mapping (described in the 2003 Afterword), but found it unsuitable for creating target domains. Mathematical functions assume the existence of the domains, seemingly ruling out the possibility of creative metaphors. We will fix that shortcoming by generating sets for morphisms. A complete analysis of the metaphorical structure of Mathematics followed in [lakoff2000mathematics], but not as a mathematical theory.
They distinguish between inherent properties and interactional properties of entities. This rhymes well with the category theoretical approach emphasizing that an object can be fully characterized by its relations to other objects (Yoneda Lemma). The overall picture of concepts cross-connected with knowledge-transferring mappings is very similar to our perspective.
2.4. Mathematical explanation and representation
At the same time, the critics also make a very typical mistake in that they necessarily mean homomorphism or even isomorphism by morphism ([pero2016varieties]), which “would not survive certain changes to the mathematical domain” ([knowles2021platonic]). We will come back to this issue and clarify the difference soon in chapter 3.
2.5. The Thousand Brains Theory of Intelligence
According to the Thousand Brains Theory ([hawkins2021thousand]), intelligence is the ability of creating predictive models of the world. The novelty of the theory lies in positing that several models of different sensory inputs exist for a single object, hence the name thousand brains, instead of having one model for one thing. Despite the multitude of models, we have stable, singular perceptions. Connections between the models and a voting mechanism ensure unambiguity.
The representations are spatial for all models, using reference frames. These are like coordinate systems but not as simple as Euclidean space. According to the hypothesis, the grid cells in the neocortex used for representing locations are reused for abstract thinking ([2019hawkinsarticle]). In this view, thinking is movement.
Our theory of understanding brings this voting mechanism for perception to a higher level of thinking. Thinking progresses from one thought to the other in a space of ideas. We find matches, morphic relations, between these spaces when we aim to understand something.
@@ 3. Morphisms: compatible operations
Here, we describe the general idea of morphisms with a degree of mathematical rigor that this discourse requires. This is a balancing act. Being too specific could make morphisms look applicable only in Mathematics. Being too general may also lead to misunderstandings. Morphisms are not just some arbitrary similarities, there are clear defining rules. For clarity, we mention upfront that
@@
  • they always go both ways (hence the need for relations);
  • and need not be totally defined (may not even require domains to exist in full).
We will talk about relations on two level: a morphism can be relational, and we can have a morphism of relations. To emphasize the dynamic nature of morphisms we introduce them in three steps, in this particular order:
@@ 3.1. Morphisms of Operations
An operation is a general term for doing something, some action. A composition of operations is just doing two operations, one after the other, let’s say $a$ after $b$. This short sequence of operations, denoted by $ab$, is another operation. It is a composite operation. However, not all operations can be connected. They should form a valid or meaningful operation together, in other words they should be composable. Two computer programs are composable if the type of the output of the first program is the same as the type of the input of the second one. Two operations can be composable one way, but not necessarily the other way. Making dough and then bake a bread form a composable pair of “kitchen operations”, but trying to bake flour, water, salt and yeast first and then trying to make a dough are not composable.
@@
We also require that composition is associative, or rather the operations we are interested in happen to be associative. This means that $(ab)c=a(bc)$, thus we can group the operations as we like, the result remains the same. Consequently, the sequence $abc$ has a well-defined meaning. It has been suggested by [wildbook] that associativity can be considered as an axiom of time.
A domain of operations is defined by a set of composable operations. We may think, if it is helpful, that the operations are of the same kind, e.g., symbol manipulations or movements in physical space. However, mathematically, we only need composability. There is no need for ontological considerations.
@@ 3.2. Morphisms of Binary Relations
3.3. Morphisms of $n$-ary relations – no composition
We can define morphisms for $n$-ary relations too. A morphism between $S$ and $T$, now domains of objects, is the function $\varphi:S\rightarrow T$, such that for all $n$-tuples $(x_{1},\ldots,x_{n})$ that are in relationship in $S$, the relation in $T$ holds for the $n$-tuple $(\varphi(x_{1}),\ldots,\varphi(x_{2}))$.
However, an $n$-ary relation in general does not have directionality like the binary case does. For $n=3$ we already have two different possibilities.
${x}$${z}$${y}$${x}$${z}$${y}$
Consequently, we do not have composition naturally. Therefore, morphisms between $n$-ary relations can only transfer static structure. This limitation also reduces the direct applicability of these morphisms.
We do not deny the usefulness of $n$-ary relations, but we claim that understanding is dynamic and compositional even for static structures. When observing an unchanging object, we have an order of looking at the parts. This sequence defines movement, which can be matched to compatible movements in other domains.
@@ 4.3. The logarithmic function
The logarithmic identities are parts of the standard curriculum, thus everyone is familiar with them.
@@ 5. Explanation, representation, understanding and communication
Understanding is a special case: an explanation given to myself. It is about bringing my thought sequence into a morphic relation with the explanandum. The condition of compatible operations means that it does not matter whether I follow my steps in thinking, only mapping the last one to the phenomenon or translating each step in my thinking and using the actions to get the same results.
@@ 5.6. The issue of (un)explainable AI
The latest contributions, such as the thought-provoking approach of [buchholz2023means], also implicitly look for some morphism, e.g. between the internal state space of the AI and some humanly interpretable domain, even by interposing some intermediate domain, e.g. by heatmapping.
@@ 6. Conclusion
We can define Mathematics as the limit of this process: using and creating morphisms in overdrive. The mathematical examples of morphisms (e.g., group and graph homomorphisms, diffeomorphisms and topological maps, functors between categories) are distilled manifestations of the mechanism of compatible operations between various domains - a necessary ingredient for any intelligence.
Much of the philosophical work ahead is to evaluate and, to the extent possible, measure the strength of morphic relations in several fields. The first candidate for morphic network analysis would be Mathematics itself: recasting the traditional approaches and describing a morphism-based philosophy of Mathematics.

### v2

@@ 1. Introduction
Finding compatible operations in two domains is not merely a tool in mathematics, but a universal mechanism for any understanding, reasoning and explanation, both for human and artificial intelligence.
These mechanisms are called morphisms (i.e., structure-preserving maps), and they could and should be used correctly and explicitly in all areas of knowledge to make philosophical problems more precise and reduce the chance of confusion arising from ambiguities.
@@
To avoid further misunderstandings, we precisely define morphisms and fix a notation in Section 2. We discuss the canonical examples in Section 3. In Section 4 we go through some explicit and implicit uses of morphisms. Section 5 elaborates on how the idea of morphisms underlie various aspects of fields related to intelligence, like explanation and understanding. We conclude in Section 6 with a summary and a research plan.
@@ 2. Morphisms: compatible operations
Here, we describe the general idea of morphisms with a degree of mathematical rigor sufficient for the current discourse. This is a balancing act. Being too specific could make morphisms look applicable only in Mathematics. Being too general may also lead to misunderstandings. Morphisms are not just some arbitrary similarities, there are clear defining rules. For clarity, we mention upfront that
@@
  • they come in pairs (we can always turn around a morphism, the reverse may cease to be a function, but it remains morphic);
  • and need not be totally defined (may not even require domains to exist in full, only given by generators).
We will talk about relations on two levels: a morphism can be relational, and we can have a morphism of relations. To emphasize the dynamic nature of morphisms we introduce them in three steps, in this particular order:
@@ 2.1. Morphisms of Operations
An operation is a general term for doing something, some action. A composition of operations is just doing two operations in sequence, one after the other. The composite, $a$ after $b$ denoted by $ab$, is another operation. We require that composition is associative, or rather the operations we are interested in happen to be associative. This means that $(ab)c=a(bc)$, thus we can group the operations as we like, the result remains the same. Consequently, the sequence $abc$ has a well-defined meaning. Associativity states that there is nothing more than the linear order of $a,b$ and $c$, thus it captures an essential property of time [wildbook].
We call composable the pairs of operations that can be connected in sequence, but not all of them can. In computing, these constraints are enforced by a type system. Two computer programs are composable if the type of the output of the first program is the same as the type of the input of the second one. Two operations can be composable one way, but not necessarily the other way. Making dough and then bake a bread form a composable pair of “kitchen operations”, but trying to bake flour, water, salt and yeast first and then trying to make a dough are not composable.
@@
A domain of operations is defined by a set of operations serving as a workspace for thinking about a given field of knowledge. This workspace is structured by composability relations, by its type system. We may think, if it is helpful, that the operations are of the same kind, e.g., symbol manipulations or movements in physical space. However, mathematically, we only need composability. There is no need for ontological considerations.
@@ 2.2. Morphisms of Binary Relations
2.3. Morphisms of $n$-ary relations – no easy composition
We can define morphisms for $n$-ary relations too. A morphism between $S$ and $T$ is the function $\varphi:S\rightarrow T$, such that for all $n$-tuples $(x_{1},\ldots,x_{n})$ that are in relationship in $S$, the relation in $T$ holds for the $n$-tuple $(\varphi(x_{1}),\ldots,\varphi(x_{2}))$.
However, an $n$-ary relation in general does not have directionality like the binary case does. We can say that the $n$ elements of related, but relation does not specify any further roles, distinction, any further connections between them. If we want to impose direction on the connections, then we have two different possibilities already for $n=3$.
${x}$${z}$${y}$${x}$${z}$${y}$${x}$${z}$${y}$
Consequently, we do not have composition naturally, not in a sequence. Therefore, morphisms between $n$-ary relations can only transfer static structure (how things are related). This limitation reduces the direct applicability of these morphisms. We do not deny the usefulness of $n$-ary relations, but we claim that understanding is dynamic even for static structures, i.e., it happens through a sequence of composable events. When observing an unchanging object, we still have an order of looking at the parts. This sequence defines movement, which can be matched to compatible movements in other domains. In a way, thinking is movement.
@@ 3.3. The logarithmic function
The logarithmic function is part of the standard curriculum, thus everyone is familiar with it. For this function, the following identity holds:
@@ 7 $W_{1}$ 8 9 $W_{2}$
4. Existing Implicit and Explicit uses of Morphisms
4.1. Category Theory
Informally, category theory can be described as an abstract theory of functions ([awodey2010category]). As a relatively concrete example of a category, we mention algebraic objects of the same kind (e.g., groups, rings), and the homomorphisms between them represented by arrows. Visually, a category is a directed graph, where associative composition is defined. For a path, a composable sequence of arrows, there is an arrow for the corresponding composition. In other words, whenever there is a way (no matter how long) from one object to another, there has to be a direct link. It is rather surprising this minimal structure can serve as the model of numerous mathematical theories ([spivak2017]).
  1. (1)
    too specific to mathematics;
  2. (2)
    too general about morphisms.
There is no contradiction here. Learning category theory requires wide and mature mathematical knowledge, although recent books ([cheng2022joy, 2022_sheaf_Rosiak]) ease this situation. Also, the exact definition of a category is geared towards algebraic objects. The need for the identity morphism for each object is less relevant for a general philosophical theory of explanation and representation. On the other hand, an abstract morphism in a category does not need to be a structure preserving map. For instance, a partially ordered set is a category. The arrows are the $\leq$ relations. These arrows merely indicate the presence of the relation without being morphisms in the literal sense.
The most relevant feature of category theory here is that it provides tools from transforming one type of mathematics into another type. It studies knowledge transferring. Moreover, category theory has a self-similar nature: a functor is a structure preserving map between categories, and natural transformation are morphisms between functors. Thus, we have a category of categories (assuming provisions against the paradoxes of naive set theory). Then, we can proceed with structure preserving maps between functors. This plurality of cross-level structure preserving maps is what we envision here for a wider context of the philosophical discourse about the interrelated notions of scientific explanation, understanding, reasoning, representation, and intelligence in general.
4.2. Algebraic Automata Theory
Algebraic automata theory studies computation with the tools of abstract algebra. We represent finite state automata as semigroups and their relationship as morphisms. One fundamental result, the prime decomposition theorem of finite state automata, states that any automaton can be built from simpler components using series-parallel connections [primedecomp65]. The control information flow is unidirectional between the components, thus the decompositions are hierarchical. An emulation, a surjective morphism from the original automaton into a nicer decomposition, allow us to understand original automaton better: we can see its subsystems and give coarse-grained descriptions of its dynamics. Decompositions are not unique, thus we can understand something in several ways.
A finite state automaton (with no specified initial and accepting states for language recognition) is a discrete dynamical system. The above decompositions introduce coordinates into these systems describing their dynamics on several levels. For instance, conserved quantities in physical systems (e.g., energy) give the top level coordinates in a hierarchical decomposition [wildbook], thus the hierarchical decomposition theory can be viewed as one basic mechanism for scientific understanding. The theory can be extended to continuous systems [KRENER1975], and a recent reformulation on categorical level [egrinagy2025sgpoid] allows further generalizations.
4.3. Cognitive Metaphors
Cognitive metaphors [lakoff2008metaphors] are thought processes where we “experience and understand one kind of things in terms of another”. Metaphor is a map between a source domain (what we know) and the target domain (what is new for us) with the purpose of conceptualizing the target. This is exactly the setup for a morphism, but the correspondence between the domains is not detailed. The authors indeed started with the idea of a mathematical mapping (described in the 2003 Afterword), but found it unsuitable for creating target domains. Mathematical functions assume the existence of the domains, seemingly ruling out the possibility of creative metaphors. This shortcoming is fixed by the generating sets for domains, we establish the morphisms for the generator elements only. It is true that metaphors have limited scope, and indeed in everyday language understanding is always partial. Chance is always there that we just ask one more question and it becomes clear that our conversion partner has different ideas. Morphisms defined for generators guarantee morphisms for the whole domain only for mathematical objects. Nevertheless, the mechanism is the same for partial agreements.
A complete analysis of the metaphorical structure of Mathematics followed in [lakoff2000mathematics], but metaphors were used only as cognitive linguistic tools external to the subject, not as mathematical objects. This was deliberate since the main purpose was to emphasize that the philosophy of Mathematics is better start with cognitive science, rather than Mathematics itself. Still, making the connection with category theory could have been beneficial.
Another similar missed opportunity in [lakoff2008metaphors] happens when the authors distinguish between inherent properties and interactional properties of entities. They argue that our concepts are primarily defined by how they interact with each other. As a perfect match, in category theory the Yoneda Lemma (see e.g.,[awodey2010category]) states that an object can be fully characterized by its relations to other objects, i.e., by the morphisms into and out of the object. Despite the omission of this connection, the cognitive metaphors’ overall picture of concepts cross-connected with knowledge-transferring mappings is very similar to our perspective.
4.4. Mathematical explanation and representation
At the same time, the critics also make a very typical mistake in that they necessarily mean homomorphism or even isomorphism by morphism ([pero2016varieties]), which “would not survive certain changes to the mathematical domain” ([knowles2021platonic]). We will come back to this issue and clarify the difference soon in chapter 2.
4.5. The Thousand Brains Theory of Intelligence
We described the process of thinking as a composable sequence of events and understanding is matching the paths through the space of ideas. This might be literally true according to the Thousand Brains Theory [hawkins2021thousand]. It defines intelligence is the ability of creating predictive models of the world. The novelty of the theory lies in positing that several models of different sensory inputs exist for a single object, hence the name thousand brains, instead of having one model for one thing. Despite the multitude of models, we have stable, singular perceptions. Connections between the models and a voting mechanism ensure unambiguity. This mechanism is like a low-level version of understanding as finding structural matches. Conversely, our theory of understanding brings this voting mechanism for perception to a higher level of thinking.
The representations are spatial for all models, using reference frames. These are like coordinate systems but not as simple as Euclidean space. According to the hypothesis, the grid cells in the neocortex used for representing locations are reused for abstract thinking [2019hawkinsarticle]. In this view, thinking is movement.
@@ 5. Explanation, representation, understanding and communication
Understanding is a special case: an explanation given to myself. It is about bringing my thought sequence into a morphic relation with the explanandum. The condition of compatible operations gives choice. I can follow steps in my thinking, only mapping the last one to the phenomenon. Or, I can translate each step in my thinking and carry out the actions or make the observations. At the end, I should get the same result.
@@ 5.6. The issue of (un)explainable AI
The latest contributions, such as the thought-provoking approach of [buchholz2023means], also implicitly look for some morphism, e.g. between the internal state space of the AI and some humanly interpretable domain, even by interposing some intermediate domain, e.g., by heatmapping.
@@ 6. Conclusion
We can define Mathematics as the limit of this process: using and creating morphisms in overdrive. The mathematical examples of morphisms (e.g., group and graph homomorphisms, diffeomorphisms and topological maps, functors between categories) are distilled manifestations of the mechanism of compatible operations between various domains - a necessary ingredient for any intelligence. However, all other fields of knowledge can benefit from the explicit description of compatible operations and their domains.